You are trying to hold your 550 g physics books with your hand against a vertical wall. Using a force sensor you know that the force you are applying is 10.0 N. Below the book on the floor you have placed a motion sensor. The position time graph for the book is given. Find the coefficient of kinetic friction.


1
Expert's answer
2021-07-15T13:13:08-0400

The book position is changing linearly with time so its velocity is constant.

this means that the net force acting on the book in the direction parallel to the surface is

"\sum F=10.0N -n =0"

Where n is the magnitude of the normal force on the book due to the wall and the net force perpendicular the surface is

"\sum F=f-0.550g=0"

Where f is the magnitude of kinetic friction and

"f=\mu n"

Where "\mu" is the coefficient of kinetic friction

Then

"n=10.0N"


"f=(0.550)g=5.39\times9.8=5.39N"

"5.39N=\mu(10.0N)\\\mu=0.539"